OpenStudy (anonymous):

Could anybody help finding the two factors of this equation: p(x)=1+(x-2)^3 one of the factor is (x-1) Thanks!

jimthompson5910 (jim_thompson5910):

1 = 1^3 you can write the right side into the form a^3 + b^3 a = 1 b = x-2 then use the sum of cubes factoring formula a^3 + b^3 = (a+b)(a^2 - ab + b^2)

OpenStudy (anonymous):

Actually I think ^3 is just for (x-2) but not 1

jimthompson5910 (jim_thompson5910):

what do you mean?

jimthompson5910 (jim_thompson5910):

p(x)=1+(x-2)^3 is the same as p(x)=1^3+(x-2)^3 since 1 = 1^3

jimthompson5910 (jim_thompson5910):

1^3 = 1*1*1 = 1

OpenStudy (anonymous):

oh true!

OpenStudy (anonymous):

omg I got it!!!! So smart! Thank you very much!

jimthompson5910 (jim_thompson5910):

you're welcome